Find the total cost of the wire needed to fence the field if the cost of fencing is $4.05 per meter. one side is 7m and the other side is 120m
Answer by jorel555(1290)

Assuming the field is rectangular, then the perimeter is
P=2l+2w=2(l+w)
(120+7)2=254
254*4.05=$1028.70 total cost for fencing!!!!!!!!!!